# CUBIC DIOPHANTINE EQUATION SATISFIABILITY IS $\Sigma^{0}_{1}$-COMPLETE

### v6

@@ Abstract
Following a mechanization in Rocq, we prove that satisfiability of single Diophantine equations of total degree $\leq 3$ over $\mathbb{N}$ (with a computably bounded number of variables depending on the instance) is $\Sigma^{0}_{1}$-complete, and hence undecidable, in the sense of Hilbert’s Tenth Problem. The core construction is a uniform primitive recursive compiler mapping arithmetic sentence codes to finite structured cubic constraint systems, such that provability in $\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1}$ is equivalent to solvability of the emitted system. Conjunction is eliminated without squaring: all emitted constraints are merged into a single cubic Diophantine equation via a two-channel equality that preserves the cubic bound without sum-of-squares. All syntactic and inference checks compile to quadratic constraints; the only cubic terms come from linear selector variables multiplying those quadratic obligations. The mechanized core is constructive and requires no classical axioms: uniform computability, checker–constraint equivalence, and the degree-$3$ bound are verified by structural induction. From this base, undecidability is obtained by a classical Diagonal Argument: any total correct decider for cubic satisfiability would compose with the compiler to decide arithmetic theoremhood.
@@ 1.1 Register Semantics
Intuition.
CircuitRegisterALUMemoryInputOutput I/O realization in Register Arithmetic ($\mathsf{RA}$) Mechanization yields mathematical programs. $\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1}$ ISA–logic duality witnesses effectivity of weak arithmetics.
Figure 1: Execution traces as finite sequences of configurations, understood arithmetically as coded finite objects, within a Von Neumann Architecture. Each step requires only Register Arithmetic $\mathsf{RA}=(\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1})\upharpoonright\mathcal{L}_{\mathsf{RA}}$.
That certain forms of Diophantine satisfiability are $\Sigma_{1}$-complete is “old news.” One obtains this by composing MRDP [Matiyasevich, 1993] with reduction techniques such as Shielding [Jones, 1980]. However, these constructions typically achieve degree bounds through auxiliary padding and simulation gadgets; the resulting systems encode computation faithfully, but the structural complexity reflects the mechanics of simulation rather than any intrinsic proof-theoretic invariant.
On the other hand, mathematical practice and computer science have accumulated a vast repertoire of methods, ranging from ad hoc constructions to fully abstract algorithms. Yet whenever any such method is carried out, it is ultimately compiled into machine code (Figure 1) and executed as a finite sequence of operations over a fixed Instruction Set Architecture (ISA).
This observation is decisive for a problem as intrinsically tied to computability as Hilbert’s 10th. Every computationally enumerable language is recognized by some finite program over a fixed instruction set; hence any reduction establishing $\Sigma_{1}$-completeness must, implicitly or explicitly, encode such a finite execution trace. In this sense, Diophantine representations inherit a structural bottleneck common to all computably enumerable sets.
Accordingly, the appropriate framework should serve simultaneously as a language of machines and of recursion, making explicit the finite operational trace underlying existential search. At its core, it should require no more than bounded arithmetic over addition—computationally no stronger than an abacus.
Definition 1.1 ($\mathsf{RA}$).
We define Register Arithmetic as a specific fragment of first-order arithmetic. The language comprises of
$\mathcal{L}_{\mathsf{RA}}\;=\;\{0,\,S,\,+,\,=\},$ (1.1)
@@
    Successor axioms reminiscent of Robinson Arithmetics,
    $S(x)\neq 0,$ (1.2)
    $S(x)=S(y)\to x=y,$ (1.3)
    $x\neq 0\to\exists y\;x=S(y),$ (1.4)
    $x+0=x,$ (1.5)
    $x+S(y)=S(x+y).$ (1.6)
@@
    $\Big(\varphi(0,\vec{y})\;\wedge\;\forall x\,\big(\varphi(x,\vec{y})\to\varphi(S(x),\vec{y})\big)\Big)\;\to\;\forall x\,\varphi(x,\vec{y}).$ (1.7)
@@
    $\forall u\;\big(\,\forall x\leq u\;\exists y\;\varphi(x,y,\vec{z})\;\to\;\exists v\;\forall x\leq u\;\exists y\leq v\;\varphi(x,y,\vec{z})\,\big).$ (1.8)
@@
$\mathsf{RA}\;=\;(\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1})\upharpoonright\mathcal{L}_{\mathsf{RA}}.$ (1.9)
Multiplication ($\mathtt{Mult}$) must not be a primitive of $\mathsf{RA}$ but it must be representable as a $\Sigma_{1}$-relation via iterated addition (Theorem 1.2). Via this representation, $\mathsf{RA}$ interprets $\mathsf{Q}$ and is interpreted in $\mathsf{EFA}$:
$\mathsf{Q}\;\leq_{\mathrm{int}}\;\mathsf{RA}+\mathtt{Mult}\;\leq_{\mathrm{int}}\;\mathsf{EFA}.$ (1.10)
The set must be also recursive, in that membership of any sentence in the axiom set is decidable. Three meta-properties of $\mathsf{RA}$ are used in the sequel:
@@
    Representability. Each primitive recursive function or relation used in this development is numeralwise representable in $\mathsf{RA}$ (Proposition 1.10).
@@
    $\operatorname{Prov}_{\mathsf{RA}}(u)\equiv\exists p\;\operatorname{Proof}_{\mathsf{RA}}(p,u)$ (1.11)
    must satisfy the Hilbert–Bernays–Löb conditions (see Theorem 3.13).
Theorem 1.2 (Representability of multiplication).
One of our first theorems is multiplication, defined as
$\mathtt{Mult}(x,y,z).$ (1.12)
@@
$t_{0}=0,\quad t_{i+1}=t_{i}+x\;\text{ for }\;i<y,\quad t_{y}=z.$ (1.14)
This relation is primitive recursive and therefore representable in $\mathsf{RA}$ (Proposition 1.10). Verification of a proposed witness trace $(t_{0},\ldots,t_{y})$ requires only $\Delta_{0}(\mathcal{L}_{\mathsf{RA}})$ checks on sequence-encoded entries, so $\mathtt{Mult}$ fits the verification floor of $\mathsf{RA}$: the theory can verify any multiplication witness but does not have multiplication as a function symbol. We treat expressions of the form $x\cdot y=z$ in $\mathsf{RA}$-internal reasoning as abbreviations for $\mathtt{Mult}(x,y,z)$. In polynomial constraints over $\mathbb{N}$ (the Diophantine object language), the symbol $\cdot$ retains its standard meaning as ring multiplication.
Intuition.
The above are implicit in the foundational theory of bounded arithmetic [Hájek and Pudlák, 1993; Buss, 1997; Cook and Nguyen, 2010]. We make their role explicit because the theory is the calibration point of this work: the weakest arithmetic in which bounded verification of finite objects—proof traces, polynomial assignments, encoded sequences—is formalizable. All internal reasoning proceeds by inspecting given encoded witnesses; no uniform totality claims are available. The theory verifies correctness of given finite witnesses; it does not certify that witnesses exist for all instances.
This distinction is visible already at the polynomial level. For any fixed assignment $\vec{x}^{*}$ and $P$ of bounded degree, the equality $P(\vec{x}^{*})=0$ is a bounded check formalizable in $\mathsf{RA}$. The existential statement $\exists\,\vec{x}\;P(\vec{x})=0$ is $\Sigma_{1}$ and lies strictly beyond such instance verification.
The cubic threshold arises naturally from this gap. All structural and syntactic verification—axiom matching, formula decoding, target equality—compiles to quadratic constraints (degree $\leq 2$), while escalation arises from enforcing bounded existential witnesses. One standard schematic guard uses
$u=1+v^{2},\quad\exists w,(\mathtt{Mult}(u,E,w)\wedge w=0),$ (1.15)
where $\mathtt{Mult}$ is the $\Sigma_{1}$-representation of multiplication by iterated addition (Theorem 1.2). This guard does not invoke multiplication as a primitive; rather, it demands a witness trace of iterated addition that $\mathtt{RA}$ can verify via $\Delta_{0}$ checks on the encoded sequence. The guard enforces that either $E=0$ or the square term forces strict positivity. Eliminating the bounded existential therefore introduces exactly one multiplicative interaction and raises the degree from $2$ to $3$. In the implemented constraint system, the actual cubic terms come from degree-$1$ selector variables multiplying quadratic obligations; the guard template is kept only as a conceptual schematic and is not multiplied by any quadratic constraint.
Lemma 1.3 (Guard correctness and degree).
@@
    Degree $4$: Jones [1980, 1982] showed that every recursively enumerable set is the solution set of a degree-44 Diophantine equation with 5858 variables. Undecidability follows classically from MRDP [Robinson et al., 1961; Matiyasevich, 1970; Davis, 1973; Matiyasevich, 1993].
@@
    The existence of a single universal cubic Diophantine equation (in a fixed number of variables) remains open; the present work establishes $\Sigma^{0}_{1}$-completeness of single cubic equations with computably bounded variable count via a direct proof-to-polynomial compiler, and attests to their constructibility.
Gasarch [2021] (cf. Definition 1.4 for notation) summarizes the state of the problem:
@@
The existence of a universal Diophantine equation of degree $3$ in a fixed number of variables sits in the open region between the fully understood quadratic case and the high-degree expressive universality constructions.
Jones explicitly singled out Degree $3$ (over $\mathbb{N}$):
@@ 1.2 Problem Semantics
Definition 1.4 ($\mathrm{H}{\mathbb{N}}(d,n)$ notation).
@@
    $P\in\mathbb{Z}[x_{1},\ldots,x_{k}]\;\text{with}\;k\leq n\;\text{and}\;\deg P\leq d,$ (1.16)
    does $P(\vec{x})=0$ have a solution in $\mathbb{N}^{k}$?” While $\mathrm{H}{\mathbb{Z}}(d,n)$ is the analogous problem with solutions in $\mathbb{Z}^{k}$.
@@
$m\in\mathbb{N}\;\Longleftrightarrow\;\exists\,x_{1},x_{2},x_{3}\in\mathbb{Z}:\;m=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{3}$ (1.17)
@@
    $\mathrm{H}{\ast}(d,n)={\rm D}$ $\;\Longrightarrow\;\mathrm{H}{\ast}(d^{\prime},n^{\prime})={\rm D},$ (1.19)
    $\mathrm{H}{\ast}(d^{\prime},n^{\prime})={\rm U}$ $\;\Longrightarrow\;\mathrm{H}{\ast}(d,n)={\rm U}.$ (1.20)
@@
    $\mathrm{H}{\mathbb{N}}(d,n)={\rm D}$ $\;\Longrightarrow\;{\rm H}\mathbb{Z}(d,n)={\rm D},$ (1.21)
    ${\rm H}\mathbb{Z}(d,n)={\rm U}$ $\;\Longrightarrow\;\mathrm{H}{\mathbb{N}}(d,n)={\rm U}.$ (1.22)
@@
    ${\rm H}\mathbb{Z}(2d,3n)={\rm D}$ $\;\Longrightarrow\;\mathrm{H}{\mathbb{N}}(d,n)={\rm D},$ (1.23)
    $\mathrm{H}{\mathbb{N}}(d,n)={\rm U}$ $\;\Longrightarrow\;{\rm H}\mathbb{Z}(2d,3n)={\rm U}.$ (1.24)
@@
By logic. Monotonicity follows since enlarging $(d,n)$ enlarges the input class, while restricting $(d,n)$ shrinks it. Domain transfer without blow-up uses sign enumeration. Difference encoding represents $x\in\mathbb{Z}$ as uvu-v with $u,v\in\mathbb{N}$, preserving degree and doubling variables. Three–square encoding replaces $x\in\mathbb{N}$ by
$x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{3},$ (1.25)
@@
Lemma 1.6 (Systems and general aggregation).
For $d,n\in\mathbb{N}$ and $\ast\in\{\mathbb{N},\mathbb{Z}\}$, let $\mathrm{H}{\ast,S}(d,n)$ denote the decision problem: “Given a finite system
$P_{1}(\vec{x})=0,\ldots,P_{m}(\vec{x})=0$ (1.26)
with $k\leq n$ variables and $\deg P_{i}\leq d$ for all $i$, determine whether there exists $\vec{x}\in\ast^{k}$ satisfying all equations.” All implications of Lemma 1.5 hold verbatim with $\mathrm{H}{\ast,S}(d,n)$ in place of $\mathrm{H}{\ast}(d,n)$. Over $\mathbb{N}$ and $\mathbb{Z}$, sum-of-squares aggregation
$P_{1}=\cdots=P_{m}=0\;\Longleftrightarrow\;\sum_{i=1}^{m}P_{i}(\vec{x})^{2}=0$ (1.27)
@@
Proposition 1.7 (Aggregation for emitted systems).
Our first contribution will be to show that valid systems emitted by the compiler of Theorem 4.31 admit an effective map producing a single equation of degree $\leq 3$ with the same solution set (up to harmless slack variables for band capacities). The construction rewrites each constraint as an equality of two nonnegative polynomials and merges all equalities into a single cubic Diophantine equation using Fibonacci-banded Zeckendorf isolation (Section 4.8); no sum-of-squares blow-up is used.

### v7

@@ Abstract
We prove that satisfiability of a single Diophantine equation of total degree $\leq 3$ over $\mathbb{N}$, with the number of variables encoded in the input instance, is $\Sigma^{0}_{1}$-complete and hence undecidable. Concretely, we give a uniform primitive recursive compiler mapping arithmetic sentence codes to cubic Diophantine constraint systems such that provability in $\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1}$ (restricted to $\mathcal{L}_{\mathsf{RA}}$) is equivalent to solvability of the emitted constraints. Degree control is local: all syntactic checks and inference verification compile to equations of degree $\leq 2$, and the only degree-$3$ terms occur when a linear selector activates a quadratic obligation. The resulting finite system is then collapsed to a single cubic equation by a degree-preserving aggregation construction.
We further internalize the compiler into a fixed four-instruction register kernel with bounded deterministic trace unrolling and a two-level aggregation layout, yielding a single explicit universal cubic polynomial in a fixed number of variables. The concrete coefficient table and a kernel-checked degree certificate are mechanized in Rocq; the semantic bridge relating arithmetic theoremhood to solvability of the pinned polynomial is proved in a separate layer without normalizing the coefficient table. Undecidability follows by diagonalization.
@@ 1.1 Register Semantics
Remark 1.1.1.
[…]
Figure 1: Execution traces as finite sequences of configurations, understood arithmetically as coded finite objects, e.g. within a Von Neumann Architecture. Each step requires only Register Arithmetic $\mathsf{RA}=(\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1})\upharpoonright\mathcal{L}_{\mathsf{RA}}$.
That certain forms of Diophantine satisfiability are $\Sigma_{1}$-complete is “old news,” trivially obtained by composing MRDP [Matiyasevich, 1993] with reduction techniques such as Shielding [Jones, 1980]. However, these constructions achieve a complexity that reflects the mechanics of simulation rather than of a proof-theoretic invariant.
Any reduction establishing $\Sigma_{1}$-completeness must encode a finite execution trace (Figure 1). The appropriate framework should therefore serve simultaneously as a language of machines and of recursion, requiring no more than bounded arithmetic over addition.
Definition 1.1.2 ($\mathsf{RA}$).
We define Register Arithmetic as a specific fragment of first-order arithmetic. The language comprises
$\mathcal{L}_{\mathsf{RA}}\;=\;\{0,\,S,\,+,\,=\},$ (1)
@@
    Successor axioms reminiscent of Robinson Arithmetic,
    $S(x)\neq 0,$ (2)
    $S(x)=S(y)\to x=y,$ (3)
    $x\neq 0\to\exists y\;x=S(y),$ (4)
    $x+0=x,$ (5)
    $x+S(y)=S(x+y).$ (6)
@@
    $\Big(\varphi(0,\vec{y})\;\wedge\;\forall x\,\big(\varphi(x,\vec{y})\to\varphi(S(x),\vec{y})\big)\Big)\;\to\;\forall x\,\varphi(x,\vec{y}).$ (7)
@@
    $\forall u\;\Big(\,\forall x\leq u\;\exists y\;\varphi(x,y,\vec{z})\;\to\;\exists v\;\forall x\leq u\;\exists y\leq v\;\varphi(x,y,\vec{z})\,\Big).$ (8)
@@
$\mathsf{RA}\;=\;(\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1})\upharpoonright\mathcal{L}_{\mathsf{RA}}.$ (9)
Multiplication ($\mathtt{Mult}$) is not a primitive of $\mathsf{RA}$ but is representable as a $\Sigma_{1}$-relation via iterated addition (Theorem 1.1.3). Via this representation, $\mathsf{RA}$ interprets $\mathsf{Q}$ and is interpreted in $\mathsf{EFA}$:
$\mathsf{Q}\;\leq_{\mathrm{int}}\;\mathsf{RA}+\mathtt{Mult}\;\leq_{\mathrm{int}}\;\mathsf{EFA}\;<\;\mathsf{PA}$ (10)
The axiom set must also be recursive, in that membership of any sentence is decidable. Three meta-properties of $\mathsf{RA}$ are used in the sequel:
@@
    Representability. Each primitive recursive function or relation used in this development is numeralwise representable in $\mathsf{RA}$ (Proposition 1.3.3).
@@
    $\operatorname{Prov}_{\mathsf{RA}}(u)\equiv\exists p\;\operatorname{Proof}_{\mathsf{RA}}(p,u)$ (11)
    must satisfy the Hilbert–Bernays–Löb conditions (see Theorem 3.5.1).
Theorem 1.1.3 (Representability of multiplication).
Our first theorem establishes the representability of multiplication, defined as
$\mathtt{Mult}(x,y,z).$ (12)
@@
$t_{0}=0,\quad t_{i+1}=t_{i}+x\;\text{ for }\;i<y,\quad t_{y}=z.$ (14)
By construction. The relation is primitive recursive and therefore representable in $\mathsf{RA}$ (Proposition 1.3.3). Verification of a proposed witness trace $(t_{0},\ldots,t_{y})$ requires only $\Delta_{0}(\mathcal{L}_{\mathsf{RA}})$ checks on sequence-encoded entries, so $\mathtt{Mult}$ fits the verification floor of $\mathsf{RA}$: the theory can verify any multiplication witness but does not have multiplication as a function symbol. We treat expressions of the form $x\cdot y=z$ in $\mathsf{RA}$-internal reasoning as abbreviations for $\mathtt{Mult}(x,y,z)$. In polynomial constraints over $\mathbb{N}$ (the Diophantine object language), the symbol $\cdot$ retains its standard meaning as ring multiplication. ∎
Remark 1.1.4.
These facts are implicit in bounded arithmetic [Hájek and Pudlák, 1993; Buss, 1997; Cook and Nguyen, 2010]. We make them explicit via the calibration point of this work: $\mathsf{RA}$ is certainly an extremely weak arithmetic in which bounded verification of finite objects is barely formalizable: it can verify correctness of given finite witnesses, yet it cannot certify witness existence for all instances.
The cubic threshold arises from this gap. Syntactic verification—axiom matching, formula decoding, target equality—compiles to quadratic constraints (degree $\leq 2$). Escalation to degree $3$ occurs when a degree-$1$ selector variable activates a quadratic obligation. In the implemented system, these selector products are the only source of cubic terms.
Lemma 1.1.5 (Guard correctness and degree).
@@
    Degree $4$: Jones [1980, 1982] showed that every recursively enumerable set is the solution set of a degree-$4$ Diophantine equation with $58$ variables. Undecidability follows classically from MRDP [Robinson et al., 1961; Matiyasevich, 1970; Davis, 1973; Matiyasevich, 1993].
@@
    The present work constructs a single universal cubic Diophantine equation in 9,6929{,}692 variables (Definition 6.1.1), establishes $\Sigma^{0}_{1}$-completeness of single cubic equations via a direct proof-to-polynomial compiler, and verifies the degree bound and universality bridge in Rocq.
Gasarch [2021] (cf. Definition 1.2.1 for notation) summarizes the state of the problem:
@@
The existence of a universal Diophantine equation of degree $3$ in a fixed number of variables was an open problem sitting between the fully understood quadratic case and the high-degree universality constructions; the present work resolves it (Section 7). Jones explicitly singled out Degree $3$ (over $\mathbb{N}$):
@@ 1.2 Problem Semantics
Definition 1.2.1 ($\mathrm{H}{\mathbb{N}}(d,n)$ notation).
@@
    $P\in\mathbb{Z}[x_{1},\ldots,x_{k}]\;\text{with}\;k\leq n\;\text{and}\;\deg P\leq d,$ (16)
    does $P(\vec{x})=0$ have a solution in $\mathbb{N}^{k}$?” The analogous problem over $\mathbb{Z}^{k}$ is denoted $\mathrm{H}{\mathbb{Z}}(d,n)$.
@@
$m\in\mathbb{N}\;\Longleftrightarrow\;\exists\,x_{1},x_{2},x_{3}\in\mathbb{Z}:\;m=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{3}$ (17)
@@
    $\mathrm{H}{\ast}(d,n)={\rm D}$ $\;\Longrightarrow\;\mathrm{H}{\ast}(d^{\prime},n^{\prime})={\rm D},$ (19)
    $\mathrm{H}{\ast}(d^{\prime},n^{\prime})={\rm U}$ $\;\Longrightarrow\;\mathrm{H}{\ast}(d,n)={\rm U}.$ (20)
@@
    $\mathrm{H}{\mathbb{N}}(d,n)={\rm D}$ $\;\Longrightarrow\;{\rm H}\mathbb{Z}(d,n)={\rm D},$ (21)
    ${\rm H}\mathbb{Z}(d,n)={\rm U}$ $\;\Longrightarrow\;\mathrm{H}{\mathbb{N}}(d,n)={\rm U}.$ (22)
@@
    ${\rm H}\mathbb{Z}(2d,3n)={\rm D}$ $\;\Longrightarrow\;\mathrm{H}{\mathbb{N}}(d,n)={\rm D},$ (23)
    $\mathrm{H}{\mathbb{N}}(d,n)={\rm U}$ $\;\Longrightarrow\;{\rm H}\mathbb{Z}(2d,3n)={\rm U}.$ (24)
@@
Immediate. Monotonicity follows since enlarging $(d,n)$ enlarges the input class, while restricting $(d,n)$ shrinks it. Domain transfer without blow-up uses sign enumeration. Difference encoding represents $x\in\mathbb{Z}$ as uvu-v with $u,v\in\mathbb{N}$, preserving degree and doubling variables. Three–square encoding replaces $x\in\mathbb{N}$ by
$x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{3},$ (25)
@@
Lemma 1.2.3 (Systems and general aggregation).
For $d,n\in\mathbb{N}$ and $\ast\in\{\mathbb{N},\mathbb{Z}\}$, let $\mathrm{H}{\ast,S}(d,n)$ denote the decision problem:
“Given a finite system
$P_{1}(\vec{x})=0,\ldots,P_{m}(\vec{x})=0$ (26)
with $k\leq n$ variables and $\deg P_{i}\leq d$ for all $i$, determine whether there exists $\vec{x}\in\ast^{k}$ satisfying all equations.”
All implications of Lemma 1.2.2 hold verbatim with $\mathrm{H}{\ast,S}(d,n)$ in place of $\mathrm{H}{\ast}(d,n)$. Over $\mathbb{N}$ and $\mathbb{Z}$, sum-of-squares aggregation
$P_{1}=\cdots=P_{m}=0\;\Longleftrightarrow\;\sum_{i=1}^{m}P_{i}(\vec{x})^{2}=0$ (27)
@@
Proposition 1.2.4 (Aggregation for emitted systems).
Our first contribution will be to show that valid systems emitted by the compiler of Theorem 4.9.4 admit an effective map producing a single equation of degree $\leq 3$ with the same solution set (up to harmless slack variables for band capacities). The construction rewrites each constraint as an equality of two nonnegative polynomials and merges all equalities into a single cubic Diophantine equation using Fibonacci-banded Zeckendorf isolation (Section 4.8); no sum-of-squares blow-up is used.